# LAPD: Langevin-Assisted Bayesian Active Learning for Physical Discovery

### v1

@@ 3.3 The Proposed Algorithms
We now introduce the procedure for implementing LAPD. This framework can be applied to scenarios with sufficient data, whether clean or noisy, utilizing reSGLD for parameter sampling, and it can also be applied to scenarios with limited data where AL will help to minimize data acquisition cost. The proposed LAPD framework is summarized in Figure 1.
@@
Figure 1: The LAPD framework is demonstrated through the identification of the Lotka-Volterra system. The process begins with Step 1: Data Preparation, where the dataset $\mathbf{U}$ is obtained from the differential equations under investigation. Step 2: Sampling: The (stochastic) gradient $\tilde{L}(\tilde{\mathbf{\Xi}}_{k})$ is computed as outlined in the Gradient Calculation procedure (bottom). Step 3: Sequential Thresholding: Based on the sampling outcome, thresholding is applied by removing bases with small absolute coefficients from the library; for example, as in the plot, the $xy$ basis for state $x$ is excluded by thresholding. If the retained library differs from the previous iteration, another sampling round is conducted. Step 4: Active Learning: If AL is adopted, continue to add new training points according to the hybrid acquisition function and restart the training. Step 5 Reconstruction: Once the final posterior distribution $P(\tilde{\mathbf{\Xi}}|\mathbf{U})$ is obtained, the system is reconstructed with uncertainty quantification.
Data preparation. To begin, prepare the time derivatives $\dot{\mathbf{U}}$ and the library $\bm{\Theta}(\mathbf{U})$. If $\dot{\mathbf{U}}$ is not directly observed, its calculation depends on the type of data. The Finite Difference in PySINDy is adopted as the derivative calculation method for clean or noisy data. For ODEs, the library is made up of polynomials of the system states. For PDEs, partial derivatives of the position $x$ will need to be calculated before constructing the library, where the finite difference is also adopted. We also attempted other approaches to approximate the derivative terms in the face of noise, like spline or Savitzky-Galoy, and found that Finite Difference best suits the proposed framework considering approximation efficiency, accuracy, and robustness.
Parameter Sampling. The posterior sampling process involves several key steps. First, select the appropriate priors for the coefficients. Next, define the posterior distribution using the negative log-posterior (6) as the energy function. To ensure that the stationary distribution of the sampler is proportional to $\exp(-L(\mathbf{\Xi}))$, set the posterior as untempered with $\tau=1$. Additionally, apply a Metropolis-Hastings (MH) adjustment to correct any bias introduced during sampling. After defining these parameters, set an appropriate burn-in period. Discard the samples from the burn-in phase and use the remaining samples to estimate the posterior distribution of the model parameters. Sequential thresholding is employed after posterior samples of the coefficients following (8), where the coefficients with posterior modes lower than some threshold $c$ are removed from the basis (Brunton et al.,, 2016), which helps to achieve sparse results.
@@
Data Acquisition for Active Learning. In the face of data scarcity, if AL is adopted, new potential design points will be added to the selected points $\mathcal{M}$. The implementation of AL in LAPD is shown in Algorithm 1. Return to Step 1 for new data preparation, for example, update the derivative and the basis library with new $\mathcal{M}$. If model training is completed, go on to reconstruct the model and estimate the uncertainty.
Model Reconstruction and Uncertainty Quantification. For ordinary differential equation (ODE) systems, predictions are made by solving the initial value problem. The solution involves integrating over the posterior distribution of $\tilde{\mathbf{\Xi}}$ using the formula:
[…]
In real applications, it is often not feasible to solve the integral in the prediction step analytically. Instead, numerical integration can be used to compute the system’s trajectory, leveraging samples of $\tilde{\mathbf{\Xi}}$ obtained from the posterior distribution. For partial differential equation (PDE) systems, given $\tilde{\mathbf{\Xi}}$, the prediction is either made by solving the system numerically, i.e., converting the system to ODE through Fast Fourier Transform (FFT), or by solving it analytically.
UQ for the coefficients comes from the samples, where the estimated distribution can be learned. UQ for the trajectory is achieved by bootstrapping from the posterior coefficient samples and make prediction with each sample.
4 Numerical Experiments with Langevin-Assisted Physical Discovery
@@
In this part, we conduct a total of four experiments: two for ODE systems (Lotka-Volterra and Lorenz) and two for PDE systems (Burgers’ Equation and Convection-Diffusion Equation). The baseline models used are SINDy (or PDE-FIND for PDEs) and UQ-SINDy. SINDy, a frequentist method, is less robust to noise compared to LAPD, as we will demonstrate later. UQ-SINDy, a Bayesian approach, provides UQ and identifies sparse systems. However, its reliance on sparsity-inducing priors can lead to failure in discovering the true system when the prior is not perfectly chosen. Additionally, we compare different Langevin MCMC methods within LAPD, including SGLD, cyclical SGMCMC, and reSGLD.
@@
Cyclical SGMCMC follows the same iteration formula as (5), but the step size is adjusted every round to balance between exploration and sampling. The stepsize at iteration $k$ is defined as
@@
where $\eta_{0}$ is the initial stepsize, $M$ is the number of cycles, and $K$ is the number of total iterations. During implementation of Langevin MCMC methods, the Metropolis-Hastings step is carried out to ensure that the distribution of the iterated instances $\mathbf{\Xi}_{k}$ sampled at round $k$ converges to the correct distribution $\Pi$ when $k\rightarrow\infty$ (Dwivedi et al.,, 2019; Chewi et al.,, 2021). Unless otherwise specified, the default sampling method adopted in this work is the 2-chain reSGLD, and the prior adopted is the regulized horseshoe with $\nu=4,s=2$ as in (Hirsh et al.,, 2022).
Before conducting the experiments, we outlined the setup details. The noise-free data of ODEs is simulated numerically with solve_ivp method from the Scipy package in Python using the ’RK45’ integration method, setting 1e-3 for relative toerance and 1e-6 for absolute tolerance. The 1d Burgers’ system was converted into a system of ODEs using FFT, and the odeint function from scipy.integrate was used to numerically generate noise free data this system. The Convection-Diffusion system was solved analytically in python with Numpy. The noisy data is generated by adding independent and identically distributed (i.i.d.) Gaussian white noise with zero mean to the clean data. The variance of the noise for each dimension of the state is calculated as $\sigma=p\cdot\text{std}(\mathbf{x})$, where $\text{std}(\mathbf{x})$ is the standard deviation of the clean data along that dimension, and $p$ is the chosen noise level (e.g., 5%). All models are trained with the same derivate $\dot{\mathbf{U}}$ and library $\mathbf{\Theta}(\mathbf{U})$ prepared, and the same threshold if sequential thresholding adopted, making sure fair comparison.
We used three error metrics to evaluate the performance of different models, namely, Error Bar, Mean Squared Error (MSE) and Akaike Information Criterion (AIC). Error Bar is uncertainty-aware and works well in the Bayesian setting. The Error Bar criterion is defined as below following Zhang and Lin, (2018):
@@
with $md$ being the total number of entries in the parameter, $\tilde{s}_{i}$ representing the estimated variance, and $\tilde{\Xi}_{i}$ denoting the estimated posterior mode. The posterior samples generated from the MCMC method allow direct calculation of the estimated $\tilde{s}_{i}$ and $\tilde{\Xi}_{i}$. If $\tilde{\Xi}_{i}=0$, this means that at location $i$, the estimated mode is equal to zero, or the estimated mode is within the threshold and therefore set to zero, then the corresponding basis would be eliminated from the library for the corresponding state. The Bayesian error bar works both in model comparison and in threshold selection. The example of threshold selection is given in Section 4.4. A lower Error Bar indicates a more reliable coefficient result. MSE quantifies the discrepancy between the predicted trajectory of the identified model and the actual measurements, which is calculated as
$MSE=\frac{1}{Nd}\sum_{i=1}^{N}\sum_{j=1}^{d}(\mathbf{U}_{ij}-\tilde{\mathbf{U}}_{ij})^{2},$
where $\mathbf{U}$ denotes the true trejectory data (noise-free) and $\tilde{\mathbf{U}}$ denotes the predicted trajectory data, $N$ and $d$ representing the number of collection and state dimension, respectively. The smaller the MSE, the better the model learned fits the data. AIC deals with the trade-off between the goodness of fit of the model and the simplicity of the model. Here, AIC can be calculated based on the MSE,
$AIC=2k+Nd\,\ln(\text{MSE}),$
where $k$ denotes the number of non-zero coefficients. The smaller the AIC, the more favorable the model, as it effectively balances model fit and parsimony, penalizing unnecessary complexity while rewarding accuracy.
@@ 4.1 Lotka-Volterra system
In this experiment, we simulate the data with the initial condition $[x_{0},y_{0}]=[10,5]$, and the system parameters $\alpha=1.0,\beta=0.1,\delta=1.5,\gamma=0.075$. Starting from $t=0$, $5,000$ time steps of states are simulated, with $\Delta t=5\cdot 10^{-3}$. The library consists of polynomial terms up to order 2, with the library basis being $[1,x(t),y(t),x(t)^{2},x(t)y(t),y(t)^{2}]$. The collected data are then corrupted with additive noise 5%.
@@
The coefficients learned from various models are presented in Table LABEL:lotka_table. The results indicate that LAPD performs well in accurately identifying the true system structure, with no extraneous or missing terms, whereas SINDy and UQ-SINDy introduce additional terms not present in the true system. LAPD also shows better performance compared to other baseline methods, as reflected in lower MSE and AIC values. UQ-SINDy, which also uses a Bayesian approach to estimate the posterior distribution of the coefficients, relies on sparsity-inducing priors such as the horseshoe prior to exclude extraneous terms. However, this method is sensitive to the choice of prior and often does not achieve the desired system sparsity.
@@
Figure 2(a) illustrates the in-sample predictions using coefficients learned from 10 different groups of priors with varying coefficients for the horseshoe prior. With $\tau_{0}\sim U(0.05,0.3)$, $\nu\sim U(0,5)$, $s\sim U(0,3)$. The plot shows considerable variability in the predictions of UQ-SINDy, with some groups failing to capture the state fluctuations. In contrast, the prediction trajectories generated by LAPD closely follow the true trajectory. This robustness is mainly due to the fact that LAPD’s coefficient updates are primarily driven by likelihood rather than the specific choice of prior, reducing sensitivity to prior selection. Additionally, the use of sequential thresholding helps achieve system sparsity rather than relying merely on sparsity-inducing priors.
@@
Figure 2(b) provides 95% confidence intervals in the prediction space. The uncertainty coverage effectively captures the true trajectory of the Lotka-Volterra system, particularly in cases where the model’s predictions deviate from the truth. When the model predictions align well with the true trajectory, the uncertainty coverage remains appropriately modest. This highlights the reliability of the LAPD in both accurate prediction and uncertainty estimation.
@@
Figure 2: (a) Predicted trajectories of UQ-SINDy and LAPD on the Lotka-Volterra system. Gray dots represent training data, purple dashed lines are trajectories generated by UQ-SINDy, and green solid lines are yielded by LAPD. Both UQ-SINDy and LAPD are run 10 times with different prior initializations. (b) Trajectory prediction and confidence interval with LAPD on Lotka-Volterra system, with the 95% confidence interval. The dashed vertical line marks the transition from in-sample prediction to out-of-sample prediction.
@@
Table 1: Learned coefficients with different models on Lotka Volterra System with 5% noise
Basis True SINDy UQ-SINDy LAPD
X Y X Y X Y X Y
1 - - 4.158e-1 - 1.814e-2 -4.512e-2 - -
x 1.000e0 - 9.858e-1 - 9.613e-1 -2.194e-2 9.966e-1 -
y - -1.500e0 - -1.428e0 1.793e-2 -1.301e0 - -1.469e0
$x^{2}$ 3.281e-4 9.838e-4 - -
$xy$1.000e-1 7.500e-2 -1.000e-1 7.446e-2 -9.707e-2 6.935e-2 -1.000e-1 7.386e-2
$y^{2}$1.957e-3 -4.288e-3 - -
Error Bar - - 1.017e3 1.747e-3
MSE - 4.621 19.541 2.926
AIC - 7.663e3 1.508e4 5.377e3

### v2

@@ 3.3 The Proposed Algorithm
We now introduce the procedure for implementing BLADE. This framework can be applied to scenarios with sufficient data, whether clean or noisy, utilizing reSGLD for parameter sampling, and it can also be applied to scenarios with limited data, where AL will help to minimize data acquisition cost. The proposed BLADE framework is summarized in Figure 1.
@@
Figure 1: The BLADE framework is demonstrated through the identification of the Lotka-Volterra system. The process begins with Step 1: Data Preparation, where the dataset $\mathbf{U}$ is obtained from the differential equations under investigation. Step 2: Sampling: The (stochastic) gradient $\tilde{L}(\tilde{\mathbf{\Xi}}_{k})$ is computed as outlined in the Gradient Calculation procedure (bottom). Step 3: Sequential Thresholding: Based on the sampling outcome, thresholding is applied by removing bases with small absolute coefficients from the library; for example, as in the plot, the $xy$ basis for state $x$ is excluded by thresholding. If the retained library differs from the previous iteration, another sampling round is conducted. Step 4: Active Learning: If AL is adopted, continue to add new training points according to the hybrid acquisition function and restart the training. Step 5 Reconstruction: Once the final posterior distribution $P(\tilde{\mathbf{\Xi}}|\mathbf{U})$ is obtained, the system is reconstructed with uncertainty quantification.
Data preparation. The first step requires constructing the time derivatives, $\dot{\mathbf{U}}$, and the library of candidate functions, $\mathbf{\Theta}(\mathbf{U})$. When $\dot{\mathbf{U}}$ is not directly measured, it is numerically approximated from the collected data $\mathbf{U}$. We employ the Finite Difference method for this task due to its computational efficiency and the low-noise characteristics of our dataset. The composition of the library $\mathbf{\Theta}(\mathbf{U})$ depends on the type of system. For ODEs, the library is composed of polynomial functions of the system’s state variables. For PDEs, the library is constructed after computing the necessary spatial partial derivatives from the data, for which the Finite Difference method is also adopted.
Parameter Sampling. Our methodology for estimating the posterior distribution of model parameters is based on a Markov Chain Monte Carlo (MCMC) sampling process. The process is initialized by defining priors for the model coefficients. The target distribution is specified by treating the negative log-posterior, $\tilde{L}(\mathbf{\Xi})$ in (6), as the energy function. To ensure the chain properly explores the target posterior, we set the temperature $\tau=1$ and incorporate a Metropolis-Hastings (MH) acceptance criterion to maintain detailed balance. An initial warm-up phase is run to guarantee the sampler has reached its stationary distribution, and these preliminary samples are discarded. The final approximation of the posterior is then constructed from all samples collected post-warm-up. Subsequently, we achieve a sparse result by applying sequential thresholding to the posterior coefficient samples from Eq.(8). In this step, any coefficient with a posterior mode smaller than a given threshold $c$ is removed from the candidate basis, as proposed in [brunton2016discovering]. After the coefficient support stabilizes, the Langevin samples quantify coefficient uncertainty within the selected model class. Accordingly, our UQ for coefficients is best interpreted as conditional on the discovered support, not as Bayesian uncertainty that marginalizes over model structure.
@@
Data Acquisition for Active Learning. In the context of data scarcity, AL is utilized to augment the set of selected points, $\mathcal{M}$, with new potential design points. The implementation of this AL procedure is formally presented in Algorithm 1. After new points are added, the data preparation step is repeated to update the derivative estimates and the basis library with the expanded set $\mathcal{M}$. Upon the completion of model training, the process advances to the reconstruction of the model and the subsequent estimation of its uncertainty.
Model Reconstruction and Uncertainty Quantification. For both ODE and PDE systems, we generate predictions and quantify uncertainty by leveraging the posterior distribution of the coefficients, $\tilde{\mathbf{\Xi}}$. For ODEs, predicting the system’s evolution requires solving the integral:
$\hat{\mathbf{x}}^{\top}(t;\tilde{\mathbf{\Xi}},\mathbf{x}_{0})=\mathbf{x}_{0}^{\top}+\int_{t_{0}}^{t}\mathbf{\Theta}(\mathbf{x}(t^{\prime}))\tilde{\mathbf{\Xi}}\,dt^{\prime}.$
As this is typically analytically intractable, we instead create an ensemble of predicted trajectories by numerically integrating the system for each posterior sample of $\tilde{\mathbf{\Xi}}$. This ensemble-based approach provides both a mean prediction and a measure of trajectory uncertainty arising from the variance in the coefficients. For PDEs, trajectories are similarly generated from each coefficient sample, using either numerical solvers (e.g., via Fast Fourier Transform) or analytical solutions where available. The uncertainty in the model coefficients themselves is directly captured by the distribution of the posterior samples.
4 System Discovery with Uncertainty Quantification
@@
To assess the performance of BLADE under conditions of ample data availability, four experiments were performed without the use of AL: two for ODE systems (Lotka-Volterra and Lorenz) and two for PDE systems (Burgers’ equation and convection–diffusion equation). The baseline models used are SINDy (or PDE-FIND for PDEs) and UQ-SINDy. SINDy, a frequentist method, is less robust to noise compared to BLADE, as we will demonstrate later. UQ-SINDy, a Bayesian approach, provides UQ and identifies sparse systems. However, its reliance on sparsity-inducing priors introduces sensitivity to prior selection, which can hinder accurate recovery of the true system when assumptions about sparsity are misaligned with the underlying dynamics. Additionally, we compare different Langevin MCMC methods within BLADE, including SGLD, cyclical SGMCMC, and reSGLD.
@@
Cyclical SGMCMC follows the same iteration formula as (5), but the step size is adjusted every round to balance between exploration and sampling. The step-size at iteration $k$ is defined as
@@
where $\eta_{0}$ is the initial step-size, $M$ is the number of cycles, and $K$ is the number of total iterations. During the implementation of Langevin MCMC methods, the Metropolis-Hastings step is performed to ensure that the distribution of the iterated instances $\tilde{\mathbf{\Xi}}_{k}$ sampled in round $k$ converges to the correct distribution $\Pi$ when $k\rightarrow\infty$ [dwivedi2019log, chewi2021optimal]. Unless otherwise specified, the default sampling method adopted in this work is the 2-chain reSGLD, and the prior adopted is the regularized horseshoe with $\nu=4,s=2$ as in [hirsh2022sparsifying]. For reSGLD, the ratio $\tilde{\sigma}^{2}/C$ in Eq. (9) is treated as a single effective hyperparameter as a common practice and is tuned during training. Empirically, good performance is obtained when the post–warm-up swapping rate between two chains stays around $5\%$$15\%$. In our experiments, we set the temperature of two chains as $\tau_{1}=1$ and $\tau_{2}=2$, and tune $\tilde{\sigma}^{2}/C$ accordingly to reach this target range.
Before conducting the experiments, we outlined the setup details. The noise-free data of ODEs is simulated numerically with solve_ivp method from the Scipy package in Python using the RK45 integration method, setting $10^{-3}$ for relative tolerance and $10^{-6}$ for absolute tolerance. The 1D Burgers’ system was converted into a system of ODEs using Fast Fourier Transform, and the odeint function from scipy.integrate was used to numerically generate noise free data for this system. The convection–diffusion system was solved analytically in Python with Numpy. The noisy data is generated by adding independent and identically distributed (i.i.d.) Gaussian white noise with zero mean is added to the clean data. The variance of the noise for each dimension of the state is calculated as $\sigma=p\cdot\text{std}(\mathbf{U}_{i})$, where $\text{std}(\mathbf{U}_{i})$ is the standard deviation of the clean data along that dimension $i$, and $p$ is the chosen noise level (e.g., 5%). All models are trained with the same derivative $\dot{\mathbf{U}}$ and library $\mathbf{\Theta}(\mathbf{U})$ prepared, and the same threshold if sequential thresholding is adopted, making sure a fair comparison. We note that the noise levels used in the ODE and PDE benchmarks are not uniform. In particular, the ODE experiments are conducted with 5% additive noise, while the PDE experiments use lower noise levels (0.2% for Burgers’ equation and 0.1% for convection–diffusion). This reflects a practical limitation of the current pipeline: for PDE discovery, numerical derivative estimation and candidate-library construction are substantially more sensitive to measurement noise. Therefore, the PDE results in this section should be interpreted as validation in low-noise regimes rather than a comprehensive stress test of high-noise robustness.
We used three error metrics to evaluate the performance of different models, namely, Total Error Bar, Mean Squared Error (MSE) and Akaike Information Criterion (AIC). Total Error Bar is uncertainty-aware and works well in the Bayesian setting. The Total Error Bar criterion is defined as below following [zhang2018robust]:
@@
with $md$ being the total number of entries in the parameter. Here, $\tilde{\Xi}_{i}$ denotes one entry of the predicted posterior mode of the coefficients and $\tilde{s}_{i}$ its posterior standard deviation; thus, Total Error Bar is a squared coefficient of variation (CV)-type aggregate uncertainty metric over the retained coefficients. Coefficients thresholded to zero are excluded to avoid numerical instability from near-zero denominators. The Bayesian Total Error Bar works both in model comparison and in threshold selection. The example of threshold selection is given in Section 4.4. A lower Total Error Bar indicates a more reliable coefficient result. MSE quantifies the discrepancy between the predicted derivatives of the identified model states and the actual measurements, which is calculated as
$\text{MSE}=\frac{1}{Nd}\sum_{i=1}^{N}\sum_{j=1}^{d}(\dot{\mathbf{U}}_{ij}-\tilde{\dot{\mathbf{U}}}_{ij})^{2},$
where $\dot{\mathbf{U}}$ denotes the true derivative (noise-free) and $\tilde{\dot{\mathbf{U}}}$ denotes the predicted derivative, $N$ and $d$ representing the number of collection and state dimension, respectively. The smaller the MSE, the better the model learned fits the data. AIC [akaike1974new, schwarz1978estimating] deals with the trade-off between the goodness of fit of the model and the simplicity of the model. Here, AIC can be calculated based on the MSE,
$\text{AIC}=2k+Nd\,\ln(\text{MSE}),$
where $k$ denotes the number of non-zero coefficients. The smaller the AIC, the more favorable the model, as it effectively balances model fit and parsimony, penalizing unnecessary complexity while rewarding accuracy. Furthermore, we assess uncertainty calibration using the Prediction Interval Coverage Probability (PICP), defined as the fraction of ground-truth values that fall within the posterior confidence interval computed from predictive samples. In addition, we report the Mean Prediction Interval Width (MPIW), defined as the average width of the same posterior confidence interval, which complements PICP by quantifying interval sharpness.
@@ 4.1 Lotka-Volterra System
In this experiment, we simulate the data with the initial condition $x_{0},y_{0}]=[10,5]$, and the system parameters $\alpha=1.0,\beta=0.1,\delta=1.5,\gamma=0.075$. Starting from $t=0$, $5,000$ time steps of states are simulated, with $\Delta t=5\cdot 10^{-3}$. The library consists of polynomial terms up to order 2, with the library basis being $1,x(t),y(t),x(t)^{2},x(t)y(t),y(t)^{2}]$. The collected data are then corrupted with 5% additive noise. For the sequential thresholding steps, the threshold is set to 0.065, which is appropriately chosen to preserve the smallest true system parameter ($\gamma=0.075$).
@@
The coefficients learned from various models are presented in Table LABEL:lotka_table. The results indicate that BLADE performs well in accurately identifying the true system structure, with no extraneous or missing terms, whereas SINDy and UQ-SINDy introduce additional terms not present in the true system. BLADE shows better performance compared to other baseline methods, as reflected in lower MSE and AIC values.
@@
Figure 2(a) illustrates the in-sample predictions using coefficients learned from 10 different groups of priors with varying coefficients for the horseshoe prior. With $\tau_{0}\sim U(0.05,0.3)$, $\nu\sim U(0,5)$, $s\sim U(0,3)$, where $U(\cdot,\cdot)$ represents uniform distribution. The plot reveals significant variability in UQ-SINDy predictions, with certain groups exhibiting discrepancies in tracking the state fluctuations. This robustness is mainly due to the fact that BLADE’s coefficient updates are primarily driven by likelihood rather than the specific choice of prior, reducing sensitivity to prior selection. Additionally, the use of sequential thresholding helps achieve system sparsity rather than relying merely on sparsity-inducing priors. UQ-SINDy, which also uses a Bayesian approach to estimate the posterior distribution of the coefficients, relies on sparsity-inducing priors such as the horseshoe prior to exclude extraneous terms. The performance of UQ-SINDy is sensitive to the choice of prior, and in some cases, may not yield the desired level of system sparsity.
@@
Figure 2(b) provides 95% confidence intervals in the prediction space. The uncertainty coverage effectively captures the true trajectory of the Lotka-Volterra system, particularly in cases where the model’s predictions deviate from the truth. When the model predictions align well with the true trajectory, the uncertainty coverage remains appropriately modest. For the Lotka–Volterra system, the 95% prediction intervals produced by the model trained on noisy data achieve a PICP of 100%. The corresponding MPIW values are 6.25 for the in-sample intervals and 16.75 for the out-of-sample intervals. Relative to the mean ground-truth prediction magnitudes of 57.85 and 60.29, respectively, these interval widths suggest that the perfect coverage is not obtained by excessively wide intervals. This highlights the reliability of the BLADE in both accurate prediction and uncertainty estimation.
@@
Figure 2: (a) Predicted trajectories of UQ-SINDy and BLADE on the Lotka-Volterra system. Gray dots represent training data, purple dashed lines are trajectories generated by UQ-SINDy, and green solid lines are yielded by BLADE. Both UQ-SINDy and BLADE run 10 times with different prior initializations. (b) Trajectory prediction and confidence interval with BLADE on the Lotka-Volterra system, with the 95% confidence interval. The dashed vertical line marks the transition from in-sample prediction to out-of-sample prediction.
@@
Table 1: Learned coefficients with different models on the Lotka-Volterra System with 5% noise
Basis True SINDy UQ-SINDy BLADE X Y X Y X Y X Y 1 - - 4.158e-1 - 1.814e-2 -4.512e-2 - - x 1.000e0 - 9.858e-1 - 9.613e-1 -2.194e-2 9.966e-1 - y - -1.500e0 - -1.428e0 1.793e-2 -1.301e0 - -1.469e0 $x^{2}$ - - - - 3.281e-4 9.838e-4 - - $xy$ -1.000e-1 7.500e-2 -1.000e-1 7.446e-2 -9.707e-2 6.935e-2 -1.000e-1 7.386e-2 $y^{2}$ - - - - -1.957e-3 -4.288e-3 - - Total Error Bar - - 1.017e3 1.747e-3 MSE - 1.248e1 4.470e1 7.977e0 AIC - 2.525e4 3.802e4 2.077e4